OpenStudy (anonymous):

What would be the steps for solving the equation 3x + 6 = 21

OpenStudy (anonymous):

subtract 6 from both sides

OpenStudy (anonymous):

And then divid by 3?

OpenStudy (anonymous):

yupp!

OpenStudy (anonymous):

And that means the answer would be 5?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

Yay! Thanks! :D

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

Subtract 6 from both sides to look like this: 3x + 6 = 21 -6 -6 now your equation is 3x = 15 x = 5

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

glad to know i helped you, but why am i blocked?

OpenStudy (anonymous):

Oops sorry about that! Not sure how that happened! Lol :)